OpenStudy (anonymous):

Which of the following equations is quadratic? x(x2 + 1) = 0 5(4x + 2) = 3 (x + 3)(x + 4) = 5

OpenStudy (anonymous):

3rd

OpenStudy (anonymous):

\[\left( x+3 \right)\left( x+4 \right) = x ^{2} + 7x +12 \] Since highest power of x is 2 So this is Quadratic Equation

OpenStudy (anonymous):

because if you multiply you get something like this x^2+7x+12

jhonyy9 (jhonyy9):

,,Corey" do you know when an equation is quadratic ?